# Robust Distributed Estimation: Extending Gossip Algorithms to Ranking and Trimmed Means

### v1

@@ Abstract
This paper addresses the problem of robust estimation in gossip algorithms over arbitrary communication graphs. Gossip algorithms are fully decentralized, relying only on local neighbor-to-neighbor communication, making them well-suited for situations where communication is constrained. A fundamental challenge in existing mean-based gossip algorithms is their vulnerability to malicious or corrupted nodes. In this paper, we show that an outlier-robust mean can be computed by globally estimating a robust statistic. More specifically, we propose a novel gossip algorithm for rank estimation, referred to as GoRank, and leverage it to design a gossip procedure dedicated to trimmed mean estimation, coined GoTrim. In addition to a detailed description of the proposed methods, a key contribution of our work is a precise convergence analysis: we establish an $\mathcal{O}(1/t)$ rate for rank estimation and an $\mathcal{O}(\log(t)/t)$ rate for trimmed mean estimation, where by $t$ is meant the number of iterations. Moreover, we provide a breakdown point analysis of GoTrim. We empirically validate our theoretical results through experiments on diverse network topologies, data distributions and contamination schemes.
@@ 1 Introduction
$\bullet$ We introduce GoTrim, the first gossip algorithm for trimmed mean estimation which can be paired with any ranking algorithm. GoTrim does not rely on strong graph topology assumptions—a common restriction in robust gossip algorithms. We prove a competitive convergence rate of order $\mathcal{O}(\log(t)/t)$ and provide a breakdown point analysis, demonstrating the robustness of its estimate.
@@ 3 A Gossip Algorithm for Distributed Ranking – GoRank
In this section, we introduce and analyze the GoRank algorithm. We establish that the expected estimates converge to the true ranks at a $\mathcal{O}(1/ct)$ rate, where the constant $c>0$ (given in Theorem 1 below) quantifies the degree of connectivity of $\mathcal{G}$: the more connected the graph, the greater this quantity and the smaller the rate bound. We also prove that the expected absolute error decreases at the same rate. In addition, we empirically validate these results with experiments involving graphs of different types, showing that the observed convergence aligns with the theoretical bounds.
@@ 3.1 Algorithm – Convergence Analysis
We state an additional result below that builds upon the previous analysis, and establishes a bound of the same order as that in Theorem 1 for the expected absolute deviation between the true ranks and the GoRank estimates.
@@
$\mathbb{E}[|R_{k}(t)-r_{k}|]\leq\frac{1}{ct}\cdot c_{n}(r_{k})\enspace,$
with $c_{n}(r_{k})=\sqrt{2n\sigma_{n}(r_{k})}$, where $c$ and $\sigma_{n}(\cdot)$ were introduced in Theorem 1.
Refer to section C in the Supplementary Material for the technical proof. Theorem 2 shows that the expected error in absolute deviation decreases at a rate of $\mathcal{O}(1/ct)$, similar to the convergence rate for the expectation $\mathbb{E}[\mathbf{R}(t)]$, but now scaled by $c_{n}(r_{k})$. In addition, Theorem 2 can be combined with Markov’s inequality to derive high-probability bounds on the ranking error. As will be shown in the next section, this is a key component in the convergence analysis of GoTrim.
@@ 4 GoTrim – A Gossip Algorithm for Trimmed Means Estimation
In this section, we present GoTrim, a gossip algorithm for trimmed means estimation. We establish a convergence in $\mathcal{O}(\log(t)/t)$ with constants depending on the data distribution and network topology.
@@ 4.1 Algorithm – Convergence Analysis
We introduce GoTrim, a gossip algorithm to estimate trimmed mean statistics (see Algorithm 2), which dynamically computes a weighted average using current estimated ranks. Notably, GoTrim can be paired with any ranking algorithm, including GoRank consequently. Let $Z_{k}(t)$ and $W_{k}(t)$ denote the local estimates of the statistic and weight at node $k$ and iteration $t$. First, by Equation 2, the $\alpha$-trimmed mean can be computed via standard gossip averaging: for all $k\in[n]$, $Z_{k}(t)=(1/n)\sum_{l=1}^{n}W_{l}(t)\cdot X_{l}$, where $W_{l}(t)=w_{n,\alpha}(R_{l}(t))$ with $w_{n,\alpha}(\cdot)$ defined in Section 2. Secondly, since ranks $R_{k}(t)$ vary over iterations, the algorithm dynamically adjusts to correct past errors: at each step, it compensates by injecting $\left(W_{k}(t)-W_{k}(t-1)\right)\cdot X_{k}$ into the averaging process. On the one hand, we have an averaging operation which is captured by the averaging matrix: $\mathbf{W}_{2}(t)=\mathbf{I}_{n}-(e_{i}-e_{j})(e_{i}-e_{j})^{\top}/2$. Similarly to the permutation matrix (see the previous section), the expectation of the averaging matrix is symmetric, doubly stochastic and has spectral gap that satisfies $10<c<$ 1 and is given by $c=\lambda_{2}/2|E|$ [31, 3]. On the other hand, we have a non-linear operation that depends on the estimated ranks: at each iteration $t>0$, each node $k$ is updated as $Z_{k}(t)=Z_{k}(t-1)+\delta_{k}(t)\cdot X_{k}$, where $\delta_{k}(t)=W_{k}(t)-W_{k}(t-1)$. Hence, the evolution of the estimates can be expressed as $\mathbf{Z}(t)=\mathbf{W}_{2}(t)\left(\mathbf{Z}(t-1)+\boldsymbol{\delta}(t)\odot\boldsymbol{X}\right),$ where $\mathbf{Z}(t)=(Z_{1}(t),\ldots,Z_{n}(t))$ and $\boldsymbol{\delta}(t)=(\delta_{1}(t),\ldots,\delta_{n}(t))$. Taking the expectation over the sampling process, the expected estimates are given by: $\mathbb{E}[\mathbf{Z}(t)]=\mathbf{W}_{2}\left(\mathbb{E}[\mathbf{Z}(t-1)]+\Delta\boldsymbol{w}(t)\odot\boldsymbol{X}\right)$ with $\mathbf{W}_{2}=\mathbf{I}_{n}-(1/2|E|)\mathbf{L}$ and $\Delta\boldsymbol{w}(t)=\mathbb{E}[\boldsymbol{\delta}(t)]$. For $t=1$, since $Z_{k}(0)=0$, we have $\mathbb{E}[\boldsymbol{Z}(1)]=\mathbf{W}_{2}\Delta\boldsymbol{w}(1)\odot\boldsymbol{X}$. Recursively, for any $t>0$,
@@
$\left|\mathbb{E}[W_{k}(t)]-w_{n,\alpha}(r_{k})\right|\leq\mathcal{O}\left(\frac{c_{n}(r_{k})}{\gamma_{k}ct}\right),$
where $\gamma_{k}=\min(\left|r_{k}-a\right|,\left|r_{k}-b\right|)\geq 1/2$ with $a=\lfloor\alpha n\rfloor+1/2$ and $b=n-\lfloor\alpha n\rfloor+1/2$ being the endpoints of interval $I_{n,\alpha}$. The constants $c$ and $c_{n}(\cdot)$ are those defined in Theorem 2.
Further details and the complete proof are provided in Section D of the Supplementary Material. Lemma 1 establishes that, for each node $k$, the estimates of the weight $W_{k}$ converge in expectation to the true weight $w_{n,\alpha}(r_{k})$ at a rate of $\mathcal{O}\left(1/\gamma_{k}ct\right)$. In addition, this lemma suggests that points closer to the interval endpoints are subject to larger errors, as reflected in the constant $\gamma_{k}$, which is consistent with the intuition that these points require greater precision in estimating ranks.
@@
Let $\mathbf{Z}(t)$ be defined as in Algorithm 2, and assume the ranking algorithm is Algorithm 1. Then, we have that for all $k\in[n]$, for any $t>T^{*}=\min\left\{t>1\,\middle|\,ct>2\log(t)\right\}$,
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\mathcal{O}\left(\frac{\log(t)}{c^{2}t}\right)\|\mathbf{K}\odot\mathbf{X}\|,$
where $\mathbf{K}=(c_{n}(r_{1})/\gamma_{1},\;\ldots,\;c_{n}(r_{n})/\gamma_{n})$ and $c$, $c_{n}(r_{k})$, $\gamma_{k}$ are the constants defined in Lemma 1.
See Section D of the Supplementary Material for the detailed proof. Theorem 3 shows that, for each node $k$, the estimate of the trimmed mean $Z_{k}$ converges in expectation to the true trimmed mean $\bar{x}_{\alpha}$ at a rate of $\mathcal{O}\left(\log(t)/c^{2}t\right)$. While the presence of the $c^{2}$ term may appear pessimistic, empirical evidence suggests that the actual convergence may be faster in practice. The vector $\mathbf{K}$ acts as a rank-dependent mask over $\mathbf{X}$, modulating the contribution of each data point $X_{k}$ to the error bound. Specifically: (1) extreme values are more heavily penalized by the mask, as they come with better rank estimation; (2) values with ranks near the trimming interval endpoints are amplified, as small inaccuracies in rank estimation can lead to disproportionately larger errors.
@@ 4.2 Robustness Analysis - Breakdown Points
$\frac{1}{n}\max\left(\left\lfloor\lfloor\alpha n\rfloor+\frac{1}{2}-\frac{K(\delta)}{t-T}\right\rfloor,0\right)\leq\varepsilon^{*}_{k}(t)\leq\frac{\lfloor\alpha n\rfloor}{n}\enspace,$
where $K(\delta)=\mathcal{O}(1/\delta)$ is a constant and $T=\mathcal{O}(\log(1/\tau\delta)$ represents the time allowed for the mean to propagate.
@@ 5 Conclusion and Future Work
We introduced and analyzed two novel gossip algorithms: GoRank for rank estimation and GoTrim for trimmed mean estimation. We proved convergence rates of $\mathcal{O}(1/t)$ and $\mathcal{O}(\log t/t)$, respectively, and established robustness guarantees for GoTrim through breakdown point analysis. Empirical results show both methods perform well on large, poorly connected networks: GoRank quickly estimates ranks, and GoTrim is robust to outliers and improves on the naive mean. Although our analysis focuses on the synchronous setting, real-world systems are often asynchronous. We present the asynchronous version of GoRank in the Supplementary Material; however, similar to Baseline++, its convergence analysis is left for future work. In addition, extending GoTrim for decentralized estimation of rank statistics [20], key tools in data analysis, e.g., for robust hypothesis testing, is another promising direction. See the Supplementary Material for further discussion.
@@ C.1 Proof of Theorem 1
$|\mathbb{E}[R_{k}(t)]-r_{i}|\leq\frac{1}{ct}\cdot\sigma_{n}\left(r_{i}\right),$ (12)
where $c=\frac{\lambda_{2}}{|E|}$, with $\lambda_{2}$ being the second smallest eigenvalue of the graph Laplacian (spectral gap), and $\sigma_{n}(r_{i})=n\tilde{\sigma}_{n}\left(\frac{r_{i}-1}{n}\right)$, with $\tilde{\sigma}_{n}(x)=\sqrt{nx(1-x)}$.
@@ C.2 Proof of Theorem 2
$\mathbb{E}[|R_{k}(t)-r_{k}|]\leq\frac{\sqrt{2}}{ct}\cdot\sqrt{n\sigma_{n}(r_{k})}\enspace,$ (18)
where $c$ is a measure of the connectivity of the network, and $\sigma_{n}(\cdot)$ is a dispersion measure.
@@
$\mathbb{E}\left[\left(R_{k}(t)-r_{k}\right)^{2}\right]\leq\left(\frac{\sqrt{2}}{ct}\right)^{2}\cdot n\sigma_{n}(r_{k}),$
@@
$R^{\prime}_{k}(t)=\frac{1}{t}\sum_{s=0}^{t-1}\mathbf{h}_{k}^{\top}\mathbf{W}_{1}(s:)\mathbf{e}_{k}.$ (19)
where for any $10\leq s\leq t-$ 1, $\mathbf{W}(s:)=\mathbf{W}(s)\ldots\mathbf{W}(0)$.
@@
$R^{\prime}_{k}(t)-r_{k}=\frac{1}{t}\sum_{s=0}^{t-1}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(s:)\mathbf{e}_{k}\enspace,$
since $\bar{h}_{k}=\mathbf{h}_{k}^{\top}\frac{\mathbf{1}_{n\times n}}{n}e_{k}$ and $\mathbf{1}_{n}^{\top}\tilde{W}_{1}(s)=0$. The squared gap can thus be expressed as follows:
$\big{(}R^{\prime}_{k}(t)-\bar{h}_{k}\big{)}^{2}=\frac{1}{t^{2}}\sum_{s=0}^{t-1}\sum_{u=0}^{t-1}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(s:)\mathbf{e}_{k}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u:)\mathbf{e}_{k}\enspace,$

### v5

@@ Abstract
This paper addresses the problem of robust estimation in gossip algorithms over arbitrary communication graphs. Gossip algorithms are fully decentralized, relying only on local neighbor-to-neighbor communication, making them well-suited for situations where communication is constrained. A fundamental challenge in existing mean-based gossip algorithms is their vulnerability to malicious or corrupted nodes. In this paper, we show that an outlier-robust mean can be computed by globally estimating a robust statistic. More specifically, we propose a novel gossip algorithm for rank estimation, referred to as GoRank, and leverage it to design a gossip procedure dedicated to trimmed mean estimation, coined GoTrim. In addition to a detailed description of the proposed methods, a key contribution of our work is a precise convergence analysis: we establish an $\mathcal{O}(1/t)$ rate for rank estimation and an $\mathcal{O}((\log t)/\smash{\sqrt{t}})$ rate for trimmed mean estimation, where by $t$ is meant the number of iterations. Moreover, we provide a breakdown point analysis of GoTrim. We empirically validate our theoretical results through experiments on diverse network topologies, data distributions and contamination schemes.
@@ 1 Introduction
$\bullet$ We introduce GoTrim, the first gossip algorithm for trimmed mean estimation which can be paired with any ranking algorithm. GoTrim does not rely on strong graph topology assumptions—a common restriction in robust gossip algorithms. We prove a competitive convergence rate of order $\mathcal{O}((\log t)/\smash{\sqrt{t}})$ and provide a breakdown point analysis, demonstrating the robustness of its estimate.
@@ 3 A Gossip Algorithm for Distributed Ranking – GoRank
In this section, we introduce and analyze the GoRank algorithm. We establish that the expected estimates converge to the true ranks at a $\mathcal{O}(1/ct)$ rate, where the constant $c>0$ (given in Theorem 1 below) quantifies the degree of connectivity of $\mathcal{G}$: the more connected the graph, the greater this quantity and the smaller the rate bound. We also prove that the expected absolute error decreases at a rate of $\mathcal{O}(1/\smash{\sqrt{ct}})$. In addition, we empirically validate these results with experiments involving graphs of different types, showing that the observed convergence aligns with the theoretical bounds.
@@ 3.1 Algorithm – Convergence Analysis
We state an additional result below that builds upon the previous analysis, and establishes a bound of order $\mathcal{O}(1/\smash{\sqrt{ct}})$ for the expected absolute deviation between the true ranks and the GoRank estimates.
@@
$\mathbb{E}[|R_{k}(t)-r_{k}|]\leq\mathcal{O}\left(\frac{1}{\sqrt{ct}}\right)\cdot\sigma_{n}(r_{k})\enspace,$
where $c$ and $\sigma_{n}(\cdot)$ were introduced in Theorem 1.
Refer to section C in the Supplementary Material for the technical proof. Theorem 2 shows that the expected error in absolute deviation decreases at a rate of $\mathcal{O}(1/\smash{\sqrt{ct}})$, similar to the convergence rate for the expectation $\mathbb{E}[\mathbf{R}(t)]$, but with a square root dependence. In addition, Theorem 2 can be combined with Markov’s inequality to derive high-probability bounds on the ranking error. As will be shown in the next section, this is a key component in the convergence analysis of GoTrim.
@@ 4 GoTrim – A Gossip Algorithm for Trimmed Means Estimation
In this section, we present GoTrim, a gossip algorithm for trimmed means estimation. We establish a convergence in $\mathcal{O}(\log(t)/\smash{\sqrt{t}})$ with constants depending on the network and data distribution.
@@ 4.1 Algorithm – Convergence Analysis
We introduce GoTrim, a gossip algorithm to estimate trimmed mean statistics (see Algorithm 2), which dynamically computes a weighted average using current estimated ranks. Notably, GoTrim can be paired with any ranking algorithm, including GoRank consequently. Let $Z_{k}(t)$ and $W_{k}(t)$ denote the local estimates of the statistic and weight at node $k$ and iteration $t$. First, by Equation 2, the $\alpha$-trimmed mean can be computed via standard gossip averaging: for all $k\in[n]$, $Z_{k}(t)=(1/n)\sum_{l=1}^{n}W_{l}(t)\cdot X_{l}$, where $W_{l}(t)=w_{n,\alpha}(R_{l}(t))$ with $w_{n,\alpha}(\cdot)$ defined in Section 2. Secondly, since ranks $R_{k}(t)$ vary over iterations, the algorithm dynamically adjusts to correct past errors: at each step, it compensates by injecting $\left(W_{k}(t)-W_{k}(t-1)\right)\cdot X_{k}$ into the averaging process. On the one hand, we have an averaging operation which is captured by the averaging matrix: $\mathbf{W}_{2}(t)=\mathbf{I}_{n}-(e_{i}-e_{j})(e_{i}-e_{j})^{\top}/2$. Similarly to the permutation matrix (see the previous section), the expectation of the averaging matrix is symmetric, doubly stochastic and has spectral gap that satisfies $10<c_{2}<$ 1 and is given by $c_{2}=c/2$ [31, 3]. On the other hand, we have a non-linear operation that depends on the estimated ranks: at each iteration $t>0$, each node $k$ is updated as $Z_{k}(t)=Z_{k}(t-1)+\delta_{k}(t)\cdot X_{k}$, where $\delta_{k}(t)=W_{k}(t)-W_{k}(t-1)$. Hence, the evolution of the estimates can be expressed as $\mathbf{Z}(t)=\mathbf{W}_{2}(t)\left(\mathbf{Z}(t-1)+\boldsymbol{\delta}(t)\odot\boldsymbol{X}\right),$ where $\mathbf{Z}(t)=(Z_{1}(t),\ldots,Z_{n}(t))$ and $\boldsymbol{\delta}(t)=(\delta_{1}(t),\ldots,\delta_{n}(t))$. Taking the expectation over the sampling process, the expected estimates are given by: $\mathbb{E}[\mathbf{Z}(t)]=\mathbf{W}_{2}\left(\mathbb{E}[\mathbf{Z}(t-1)]+\Delta\boldsymbol{w}(t)\odot\boldsymbol{X}\right)$ with $\mathbf{W}_{2}=\mathbf{I}_{n}-(1/2|E|)\mathbf{L}$ and $\Delta\boldsymbol{w}(t)=\mathbb{E}[\boldsymbol{\delta}(t)]$. For $t=1$, since $Z_{k}(0)=0$, we have $\mathbb{E}[\boldsymbol{Z}(1)]=\mathbf{W}_{2}\Delta\boldsymbol{w}(1)\odot\boldsymbol{X}$. Recursively, for any $t>0$,
@@
$\left|\mathbb{E}[W_{k}(t)]-w_{n,\alpha}(r_{k})\right|\leq\mathcal{O}\left(\frac{1}{\gamma_{k}\sqrt{ct}}\right)\cdot\sigma_{n}(r_{k}),$
where $\gamma_{k}=\min(\left|r_{k}-a\right|,\left|r_{k}-b\right|)\geq 1/2$ with $a=\lfloor\alpha n\rfloor+1/2$ and $b=n-\lfloor\alpha n\rfloor+1/2$ being the endpoints of interval $I_{n,\alpha}$. The constants $c$ and $\sigma_{n}(\cdot)$ are those defined in Theorem 2.
Further details and the complete proof are provided in Section D of the Supplementary Material. Lemma 1 establishes that, for each node $k$, the estimates of the weight $W_{k}$ converge in expectation to the true weight $w_{n,\alpha}(r_{k})$ at a rate of $\mathcal{O}\left(1/\gamma_{k}\smash{\sqrt{ct}}\right)$. In addition, this lemma suggests that points closer to the interval endpoints are subject to larger errors, as reflected in the constant $\gamma_{k}$, which is consistent with the intuition that these points require greater precision in estimating ranks.
@@
Let $\mathbf{Z}(t)$ be defined as in Algorithm 2, and assume the ranking algorithm is Algorithm 1. Then, we have that for all $k\in[n]$, for any $t>T^{*}=\min\left\{t>1\,\middle|\,ct>\log(tc^{-1})\right\}$,
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\mathcal{O}\left(\frac{\log(tc^{-1})}{\sqrt{c^{2}t}}\right)\|\mathbf{K}\odot\mathbf{X}\|,$
where $\mathbf{K}=(\sigma_{n}(r_{1})/\gamma_{1},\;\ldots,\;\sigma_{n}(r_{n})/\gamma_{n})$ and $c$, $\sigma_{n}(r_{k})$, $\gamma_{k}$ are the constants defined in Lemma 1.
See Section D of the Supplementary Material for the detailed proof. Theorem 3 shows that, for each node $k$, the estimate of the trimmed mean $Z_{k}$ converges in expectation to the true trimmed mean $\bar{x}_{\alpha}$ at a rate of $\mathcal{O}\left(\log(tc^{-1})/\smash{\sqrt{c^{2}t}}\right)$. While the presence of the $c^{2}$ term may appear pessimistic, empirical evidence suggests that the actual convergence may be faster in practice. The vector $\mathbf{K}$ acts as a rank-dependent mask over $\mathbf{X}$, modulating the contribution of each data point $X_{k}$ to the error bound. Specifically: (1) extreme values are more heavily penalized by the mask, as they come with better rank estimation; (2) values with ranks near the trimming interval endpoints are amplified, as small inaccuracies in rank estimation can lead to disproportionately larger errors.
@@ 4.2 Robustness Analysis - Breakdown Points
$\frac{1}{n}\max\left(\left\lfloor\lfloor\alpha n\rfloor+\frac{1}{2}-\frac{K(\delta)}{\sqrt{t-T}}\right\rfloor,0\right)\leq\varepsilon^{*}_{k}(t)\leq\frac{\lfloor\alpha n\rfloor}{n}\enspace,$
where $K(\delta)=\mathcal{O}(1/\delta)$ is a constant and $T=\mathcal{O}(\log(1/\tau\delta))$ represents the time allowed for the mean to propagate.
@@ 5 Conclusion and Future Work
We introduced and analyzed two novel gossip algorithms: GoRank for rank estimation and GoTrim for trimmed mean estimation. We proved convergence rates of $\mathcal{O}(1/t)$ and $\mathcal{O}((\log t)/\smash{\sqrt{t}})$, respectively, and established robustness guarantees for GoTrim through breakdown point analysis. Empirical results show both methods perform well on large, poorly connected networks: GoRank quickly estimates ranks, and GoTrim is robust to outliers and improves on the naive mean. Although our analysis focuses on the synchronous setting, real-world systems are often asynchronous. We present the asynchronous version of GoRank in the Supplementary Material; however, similar to Baseline++, its convergence analysis is left for future work. In addition, extending GoTrim for decentralized estimation of rank statistics [20], key tools in data analysis, e.g., for robust hypothesis testing, is another promising direction. See the Supplementary Material for further discussion.
@@ C.1 Proof of Theorem 1
where $c=\frac{\lambda_{2}}{|E|}$, with $\lambda_{2}$ being the second smallest eigenvalue of the graph Laplacian (spectral gap), and $\sigma_{n}(r_{k})=n\tilde{\sigma}_{n}\left(\frac{r_{k}-1}{n}\right)$, with $\tilde{\sigma}_{n}(x)=\sqrt{nx(1-x)}$.
@@ C.2 Proof of Theorem 2
$\mathbb{E}[|R_{k}(t)-r_{k}|]\leq\sqrt{\frac{3}{ct}}\cdot\sigma_{n}(r_{k})\enspace,$ (18)
where $c$ and $\sigma_{n}(\cdot)$ are defined in Theorem 1.
@@
We have
$\mathbb{E}\left[\left(R_{k}(t)-r_{k}\right)^{2}\right]\leq\frac{3}{ct}\cdot\sigma_{n}\left(r_{k}\right)^{2},$
@@
$R^{\prime}_{k}(t)=\frac{1}{t}\sum_{s=0}^{t-1}\mathbf{h}_{k}^{\top}\mathbf{W}_{1}(s:)^{\top}\mathbf{e}_{k}.$ (19)
where for any $11\leq s\leq t-$ 1, $\mathbf{W}(s:)=\mathbf{W}(s)\ldots\mathbf{W}(1)$ and $\mathbf{W}(0:)=\mathbf{I}_{n}$.
@@
$R^{\prime}_{k}(t)-r_{k}=\frac{1}{t}\sum_{s=0}^{t-1}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}^{\top}_{1}(s:)\mathbf{e}_{k}\enspace,$
since $\bar{h}_{k}=\mathbf{h}_{k}^{\top}\frac{\mathbf{1}_{n\times n}}{n}e_{k}$ and $\mathbf{1}_{n}^{\top}\mathbf{\tilde{W}}_{1}(s)=0$. The squared gap can thus be expressed as follows:
$\big{(}R^{\prime}_{k}(t)-\bar{h}_{k}\big{)}^{2}=\frac{1}{t^{2}}\sum_{s=0}^{t-1}\sum_{u=0}^{t-1}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(s{:})^{\top}\mathbf{e}_{k}(\mathbf{h}_{k}-\bar{h}_{k}\mathbf{1}_{n})^{\top}\tilde{\mathbf{W}}_{1}(u{:})^{\top}\mathbf{e}_{k}\enspace,$